OpenStudy (anonymous):

can you help me solve this equation? -7=h over 5

OpenStudy (anonymous):

\[-7=\frac{h}{5}\iff h=-7\times 5\]

OpenStudy (anonymous):

i do not understand this at all??? :(

OpenStudy (anonymous):

multiply both sides by 5? as he did.

OpenStudy (anonymous):

i got a number, i don't tell you what it is, but i tell you if you divide the number by \(7\) you get \(-5\)

OpenStudy (anonymous):

it is pretty obvious that that number has to be \(-35\) since \(-35\div 5=-7\)

OpenStudy (anonymous):

which is another way of asking, what is \(-7\times 5\)

OpenStudy (anonymous):

ohhhh okay okay okay i see what you did there!!

OpenStudy (anonymous):

thank you ive been showed this over and over and it just clicked thanks =]

OpenStudy (skullpatrol):

-7 equals some number divided by 5 What is the number?

OpenStudy (anonymous):

good i think i kind of said it wrong but the idea was right in the equation \[-7=\frac{h}{5}\] you only have to numbers to work with \(-7\) and \(5\) it is really the same as asking what is \(-7\times 5\)

OpenStudy (anonymous):

-35 right?

OpenStudy (anonymous):

right

OpenStudy (skullpatrol):

-7 equals -35 divided by 5 <-----True :D